# Monochromatic Translated Product and Answering Sahasrabudhe’s Conjecture

### v2

@@ Abstract
This article resolves two related problems in Ramsey theory on the integers. We show that for any finite coloring of the set of natural numbers, there exist numbers $a$ and $b$ for which the configuration $\{a,b,ab,(a+1)b\}$ is monochromatic. This finding has two main consequences: first, it disproves a conjecture proposed by J. Sahasrabudhe; second, it establishes a variant of the long-standing Hindman’s conjecture, which asks for a monochromatic set of the form $\{a,b,ab,a+b\}$.
Mathematics subject classification 2020: 05D10.
Keywords: Schur Theorem, IPIP sets, Hindman Conjecture, Moreira’s sum product theorem.
@@ 1 Introduction
For more than a century, it remained unknown whether these two phenomena could be unified in a single statement. In 1979, R. L. Graham and N. Hindman independently established such a result for two-colorings (see [2, pp. 68–69], [4, Section 4]). Subsequently, N. Hindman conjectured that the statement should hold for all finite colorings, a problem now known as the Hindman conjecture. The strongest known partial result toward this conjecture is due to J. Moreira [6], who proved that there exist $a,b\in\mathbb{N}$ such that the set $\{a,ab,a+b\}$ is monochromatic; see also [1] for a concise proof.
In 2017, while investigating asymmetric variants of Schur’s theorem, J. Sahasrabudhe ( see [7, Question 31.]) conjectured that there exist finite colorings of the natural numbers that avoid monochromatic configurations of the form $\{a,b,b(a+1)\}$. In this paper, we disprove this conjecture by showing that every finite coloring of $\mathbb{N}$ necessarily contains such a configuration. In fact, we establish a stronger result: the larger set $\{a,b,ab,a(b+1)\}$ can always be found monochromatic.
Our result can therefore be viewed as a variant of the Hindman conjecture. The technique of our proof builds upon the argument of Moreira [6], who employed van der Waerden’s theorem, whereas we make iterative use of Hindman’s theorem.
@@
Theorem $1.$
For any finite coloring of $\mathbb{N}$, there exist $a,b\in\mathbb{N}$ such that the set $\{a,b,ab,(a+1)b\}$ is monochromatic.
2 Proof of Theorem 1.1
Before we proceed to the proof of our Theorem 1.1, we need to recall some technical definitions.
Definition 2.1.
@@
Lemma 2.3.
For any IP set $A\subseteq\mathbb{N}$ and any natural number $y$, there exists an another IP set $A^{\prime}\subseteq A$ satisfying $y^{-1}A^{\prime}\subseteq\mathbb{N}$.
@@
It suffices to show that for every $y\in\mathbb{N}$, there exists an $R\in\mathbb{N}$ such that any $IP_{R}$ (infinitely partition regular) set contains an element divisible by $y$. Once this is established, we may partition the original IP set into disjoint $IP_{R}$ subsets and select one element from each of them.
By Hindman’s theorem and a compactness argument, there exists an $R\in\mathbb{N}$ such that any $y$-coloring of an $IP_{R}$ set contains a monochromatic $\operatorname{IP}_{y}$ subset. Given an $IP_{R}$ set AA, define the coloring
$c:A\to\{0,1,\ldots,y-1\},\qquad c(a)=a\bmod y.$
By the choice of $R$, there exist $y$ elements of AA having the same residue modulo $y$; hence their sum belongs to AA and is divisible by $y$. This completes the proof. ∎
Lemma 2.4.
@@
$A=\left\{\sum_{n\in F}x_{n}:\varnothing\neq F\in\mathcal{P}_{f}(\mathbb{N})\right\}.$
@@
Now we are in the position to prove our Theorem $1.$
Proof of Theorem 1.1:.
@@
Without loss of generality, assume that $A_{1}$ is an IPIP set, and set $D_{0}=A_{1}$. By Lemma 2.4, we can choose an element $y_{1}\in A_{1}$ such that $(-y_{1}+A_{1})\cap A_{1}$ is an IPIP set. Then, by Lemma 2.3, we define
$C_{1}=y_{1}^{-1}\big((-y_{1}+A_{1})\cap A_{1}\big),$
which is a subset of $\mathbb{N}$. Finally, by Lemma 2.2, there exists $i_{1}\in\{1,2,\ldots,r\}$ such that
@@
Using Lemmas 2.3 and 2.4, choose $y_{2}\in D_{1}$ such that
$C_{2}=y_{2}^{-1}\big((-y_{2}+D_{1})\cap D_{1}\big)\subset\mathbb{N}$
is an IPIP set. Then, by Lemma 2.2, there exists $i_{2}\in\{1,2,\ldots,r\}$ such that $D_{2}=C_{2}\cap A_{i_{2}}$ is an IPIP set. Proceeding inductively, we define for each $n\in\mathbb{N}$:
  1. $(1)$
    choose $y_{n}\in D_{n-1}$ and $i_{n}\in\{1,2,\ldots,r\}$ such that,
  2. $(2)$
    the set
@@ 2. $(2)$
$xy_{n}\cdots y_{j+1}=ab\in A_{k},\quad\text{and}\quad(xy_{n}+y_{n})\,y_{n-1}\cdots y_{j+1}=(a+1)b\in A_{k}.$
@@
Acknowledgment
The author is thankful to the referees for their comments on the previous draft of this paper and also for the proof of the Lemma 2.3. This paper is supported by NBHM postdoctoral fellowship with reference no: 0204/27/(27)/2023/R & D-II/11927.

### v3

@@ Abstract
This article resolves two related problems in Ramsey theory on the integers. We show that for any finite coloring of the set of natural numbers, there exist numbers $a$ and $b$ for which the configuration $\{a,b,ab,a(b+1)\}$ is monochromatic. By redefining the variables $a=x$ and $ab=y,$ our configurations transforms into $\{x,y,x+y,\frac{y}{x}\}.$ This finding has two main consequences: first, it disproves a conjecture proposed by J. Sahasrabudhe; second, it establishes a quotient version of the long-standing Hindman’s conjecture, which asks for a monochromatic set of the form $\{x,y,x+y,xy\}$.
[…]
[classification=text]
@@ 1 Introduction
For more than a century, it remained unknown whether these two phenomena could be unified in a single statement. That is, whether $\{a,b,a+b,a\cdot b\}$ is also partition regular (see [2]). In 1979, R. L. Graham and N. Hindman independently established such a result for two-colorings (see [4, Section 4]). Subsequently, N. Hindman conjectured that the statement should hold for all finite colorings, a problem now known as the Hindman conjecture. The strongest known partial result toward this conjecture is due to J. Moreira [6], who proved that there exist $a,b\in\mathbb{N}$ such that the set $\{a,ab,a+b\}$ is monochromatic; see also [1] for a concise proof.
In 2017, while investigating asymmetric variants of Schur’s theorem, J. Sahasrabudhe ( see [7, Question 31.]) conjectured that there exist finite colorings of the natural numbers that avoid monochromatic configurations of the form $\{a,b,a(b+1)\}$. In this paper, we disprove this conjecture by showing that every finite coloring of $\mathbb{N}$ necessarily contains such a configuration. In fact, we establish a stronger result: the larger set $\{a,b,ab,a(b+1)\}$ can always be found monochromatic.
On the other hand, by redefining the variables $a=x,ab=y$ our result shows that the pattern $\{x,y,x+y,\frac{y}{x}\}$ is monochromatic. Our result can therefore be viewed as a quotient version of the Hindman conjecture. The technique of our proof builds upon the argument of Moreira [6], who employed van der Waerden’s theorem, whereas we make iterative use of Hindman’s theorem.
@@
Theorem 1.
For any finite coloring of $\mathbb{N}$, there exist $a,b\in\mathbb{N}$ such that the set $\{a,b,ab,a(b+1)\}$ is monochromatic.
2 Proof of Theorem 1
Before we proceed to the proof of our Theorem 1, we need to recall some technical definitions.
Definition 1.
@@ 2 Proof of Theorem 1
The proof of the next two lemmas is standard, but we give the proofs here for the sake of completeness.
Lemma 2.
For any IP set $A\subseteq\mathbb{N}$ and any natural number $y$, there exists another IP set $A^{\prime}\subseteq A$ satisfying $y^{-1}A^{\prime}\subseteq\mathbb{N}$.
@@
By passing to a subsequence if necessary, we may assume that AA contains $FS(\langle x_{n}\rangle)$ for some sequence $\langle x_{n}\rangle\subset(y\mathbb{N}+j)$, where $j\in\{0,1,\ldots,y-1\}$.
Now let $F$ be any finite set with $|F|=y$. Then
$x_{F}=\sum_{n\in F}x_{n}\in y\mathbb{N}.$
Thus every sum of exactly $y$ distinct terms from the sequence is divisible by $y$.
Selecting infinitely many distinct elements from the sequence $\langle x_{F}\rangle$ and using them as generators, it follows that AA contains another finite sums set
$A^{\prime}=FS(\langle y_{n}\rangle),$
where each $y_{n}$ is of the form
$y_{n}=yz_{n}$
for some integer $z_{n}$. And hence
$FS(\langle z_{n}\rangle)=y^{-1}A^{\prime}\subseteq\mathbb{N}.$
This completes the proof. ∎
Lemma 3.
@@
$A=\left\{\sum_{n\in F}x_{n}:F\in\mathcal{P}_{f}(\mathbb{N})\right\}.$
@@
Now we are in the position to prove our Theorem 1.
Proof of Theorem 1:.
@@
Without loss of generality, assume that $A_{1}$ is an IPIP set, and set $D_{0}=A_{1}$. By Lemma 3, we can choose an element $y_{1}\in A_{1}$ such that $(-y_{1}+A_{1})\cap A_{1}$ is an IPIP set. Then, by Lemma 2, we can choose an IPIP set
$C_{1}\subset y_{1}^{-1}\big((-y_{1}+A_{1})\cap A_{1}\big),$
which is a subset of $\mathbb{N}$. Finally, by Lemma 1, there exists $i_{1}\in\{1,2,\ldots,r\}$ such that
@@
Using Lemmas 2 and 3, choose $y_{2}\in D_{1}$ such that
$C_{2}\subseteq y_{2}^{-1}\big((-y_{2}+D_{1})\cap D_{1}\big)\subset\mathbb{N}$
is an IPIP set. Then, by Lemma 1, there exists $i_{2}\in\{1,2,\ldots,r\}$ such that $D_{2}=C_{2}\cap A_{i_{2}}$ is an IPIP set. Proceeding inductively, for each $n\in\mathbb{N}$, choose $y_{n}\in D_{n-1}$ and $i_{n}\in\{1,2,\ldots,r\}$ such that the set
@@
$xy_{n}\cdots y_{j+1}=ab\in A_{k},\quad\text{and}\quad(xy_{n}+y_{n})\,y_{n-1}\cdots y_{j+1}=a(b+1)\in A_{k}.$
@@
Acknowledgments
The author is thankful to the referees for their comments on the previous draft of this paper and also for the proof of the Lemma 2. This paper is supported by NBHM postdoctoral fellowship with reference no: 0204/27/(27)/2023/R & D-II/11927.